OpenStudy (anonymous):

. A new medical test provides a false positive result for Hepatitis 2% of the time. That is, a perfectly healthy subject being tested for Hepatitis will test as being infected 2% of the time. In research, the test is given to 30 healthy (not having Hepatitis) subjects. Let X be the number of subjects who test positive for the disease. a) What is the probability that all 30 subjects will appropriately test as not being infected? b) What are the mean and standard deviation of X? c) To what extent do you think this is a viable test to use in the field of medicine? (3 points)

OpenStudy (kropot72):

The probability of a true positive is 1.00 - 0.02 = 0.98 The probability that 30 healthy subjects will test as not being infected is given by: \[(0.98)^{30}=you\ can\ calculate\]

OpenStudy (anonymous):

(0.98)^30=29.4

OpenStudy (kropot72):

Not really. You have multiplied 0.98 by 30, however (0.98)^30 means 0.98 multiplied by itself 30 times.

OpenStudy (anonymous):

not understanding what you are saying

OpenStudy (anonymous):

0.54548

OpenStudy (kropot72):

For example: \[(0.98)^{5}=0.98\times0.98\times0.98\times0.98\times0.98=0.90392\]

OpenStudy (kropot72):

You are correct in calculating the required probability to be 0.54548.

OpenStudy (anonymous):

where did (0.98)^5 came from

OpenStudy (kropot72):

That was an example to show you the meaning of raising 0.98 to a power (the power of 5 as an example). Of course solving the problem requires 0.98 to be multiplied by itself the appropriate number of times which you did to get the correct answer of 0.54548.

OpenStudy (anonymous):

ok at first I put it the calculator wrong

OpenStudy (anonymous):

so a) is 0.54548

OpenStudy (anonymous):

into calculator

OpenStudy (kropot72):

(b) The mean is given by: \[np=30\times0.98=you\ can\ calculate\] The standard deviation is given by: \[\sqrt{np(1-p)}=\sqrt{30\times0.98\times0.02}=you\ can\ calculate\]

OpenStudy (anonymous):

mean= 29.4 and sd=0.10

OpenStudy (anonymous):

or is it 0.11

OpenStudy (anonymous):

I think it is a viable test use in the field of medicine because of the result from the probablility.

OpenStudy (kropot72):

The s.d. is 0.7668.

OpenStudy (anonymous):

mean is 6

OpenStudy (kropot72):

The mean of X is np = 30 * 0.98 = 29.4